A company currently pays a dividend of $3.5 per share (D0 = $3.5). It is estimated that the company's dividend will grow at a rate of 25% per year for the next 2 years, and then at a constant rate of 5% thereafter. The company's stock has a beta of 0.9, the risk-free rate is 3%, and the market risk premium is 6%. What is your estimate of the stock's current price? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Required Return, rs = Risk-free Rate + Beta * Market Risk Premium
Required Return, rs = 3.00% + 0.90 * 6.00%
Required Return, rs = 8.40%

Last Dividend, D0 = $3.50

Growth rate for next 2 years is 25% and a constant growth rate (g) of 5% thereafter

D1 = $3.50 * 1.25 = $4.375
D2 = $4.375 * 1.25 = $5.46875
D3 = $5.46875 * 1.05 = $5.7421875

P2 = D3 / (rs - g)
P2 = $5.7421875 / (0.084 - 0.05)
P2 = $168.8879

P0 = $4.375/1.084 + $5.46875/1.084^2 + $168.8879/1.084^2
P0 = $152.42

So, the current stock price is $152.42